OpenStudy (anonymous):

Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x².

OpenStudy (bibby):

Remember that taking the square root of a number is raising said number to the 1/2 power. As Such \[(\sqrt{x})^{2} = x || (\sqrt{x})^{4} = x ^{2} || (\sqrt[3]{x})^6 = x ^{2}\]